OpenStudy (ebonyserrano):

Three consecutive odd integers have a sum of 45. Find the integers.

OpenStudy (ebonyserrano):

i think it might be 15,16, and 17

OpenStudy (derricklewis):

It is 15,16,17 plzz give me a medal

OpenStudy (mrnood):

16 is not odd they do not add up to 45 apart from that - you deserve a medal

OpenStudy (mrnood):

@Ebonyserrano if the middle of the 3 numbers is n (and is ODD) how would oyu write an expression for the NEXT ODD number??

OpenStudy (holsteremission):

What was wrong with the other two times you posted this question? Both questions got answered. Seems to me you're not actually willing to do the work.

OpenStudy (derricklewis):

X+X+2+X+4=45 3X+6=45 3X=45-6 3X=39

OpenStudy (derricklewis):

hOPE I HELP ED

OpenStudy (mrnood):

@derricklewis so how did you help by confirming his WRONG answer (and demanding a medal?)

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (will.h):

Lol MrNood

OpenStudy (will.h):

Anyways. You'll need to make an expression

OpenStudy (will.h):

X + (x + 2) + (x + 4) = 45 Solve for x 1st 3x + 6 = 45 3x = 39 X = 13 Now substitute that in the expression above The 1st integer would be 13 The 2nd would be 15 The 3rd would be 17 Let's see it they add up to 45 13 + 15 + 17 = 45 45 = 45 Therefore the assumption was confirmed

OpenStudy (will.h):

Make sense?

OpenStudy (ebonyserrano):

yea thank you